OpenStudy (anonymous):

Simplify: 2x(x + 18) - x(4x - 2) >> -2x2 + 38x -2x2 + 38 -2x2 + 34 I think its the first one, BUT isn't the first one and the second they same(yes i put them on here right)

OpenStudy (goldphenoix):

2x(x+18) = 2x^2 + 36x -x(4x-2) = -4x^2 + 2x 2x^2 - 4x^2 = -2x^2 36x + 2x = 38x. -2x^2 + 38x I also agree with the first one.

OpenStudy (anonymous):

I would distribute then factor. BUt it could 2 much steps

hero (hero):

2x(x + 18) - x(4x - 2) 2x(x + 18) -2x(2x - 1) 2x(x + 18 - (2x - 1)) 2x(x + 18 - 2x + 1) 2x(x - 2x + 18 + 1) 2x(-x + 19) -2x^2 + 38x